# The Nonlinear Filter Model of Stream Cipher Redivivus

### v1

@@ 4 Concrete Proposals
Given $\kappa$, we obtained representative values of $L$ and $m$. The procedure we followed is to choose the value of $L$ to be the first prime number greater than $2\kappa$, and then for the chosen value of $L$, choose $m$ to be the least integer such that $\beta,\gamma>2^{\kappa}$. This ensures security against algebraic attacks considered above. Next, using $\kappa$ and the corresponding values of $L$ and $m$, we computed the maximum value of $B$ such that $B\leq\kappa$, $\alpha>2^{B}$ and (9), (10) and (11) hold. In each of the cases that we considered, it turns out that this maximum value of $B$ is in fact $\kappa$. So for the chosen values of $L$ and $m$, the corresponding $\mathcal{S}(L,m)$ ensures $\kappa$-bit security against the above analysed correlation attacks even when the adversary has access to $2^{\kappa}$ keystream bits. We note, however, that generating $2^{\kappa}$ keystream bits from a single key and IV pair is meaningless from a practical point of view. Instead, we set $B=64$, i.e. from a single key and IV pair at most $2^{64}$ keystream bits are to be generated. Table 1 shows the values of $\kappa$, $L$, $m$ and other parameters of the filtering function. We note the following points regarding the entries in Table 1.
@@
    The number of variables of the filtering function is $12m+$ 1 (and not $m$).
@@
    The feedback polynomial for the LFSR has to be a primitive polynomial of degree $L$. The above analysis of correlation attacks shows that for the obtained values of $L$ and $m$, it does not matter whether the polynomial is sparse, or whether it has a sparse multiple. So from an efficiency point of view, one may choose a primitive polynomial of the least possible weight, possibly a trinomial or a pentanomial.
@@ 5 Efficiency of Computing ${\sf MM}_{n}$
In Section 4, we proposed using ${\sf MM}_{2m+1}$ as the filtering function in the filter model. Further, Table 1 provides specific suggestions of values of $m$ for achieving different security levels. In this section, we consider the complexity of implementing ${\sf MM}_{2m+1}$. Since ${\sf MM}_{2m+1}$ is constructed from ${\sf MM}_{2m}$ using one XOR gate, we need to consider the complexity of implementing ${\sf MM}_{2m}$.
@@
A half-adder takes two input bits and outputs two bits which represent sum of the two input bits. A full adder takes three input bits and outputs two bits which represent the sum of the three input bits. We estimate the numbers of half and full adders that are required for computing the weight of an $m$-bit string $\mathbf{x}$. The algorithm for computing weight that we use is from [6]. If $m=2$, a half adder computes the weight, and if $m=3$, a full adder computes the weight. For $m>3$, write $m=m_{1}+m_{2}+1$, where $m_{1}+1$ is the highest power of two that is at most $m$. The algorithm computes the weight of the first $m_{1}$ bits of $\mathbf{x}$, the weight of the next $m_{2}$ bits of $\mathbf{x}$, and then adds these two weights together together with the last bit of $\mathbf{x}$.
Using the above algorithm, it is easy to show that the computation of the weight of an $m$-bit string, when $m=2^{r}-1$ with $r\geq 1$, requires $12^{r}-r-$ 1 full adders. From this it easily follows that the number of full adders required to compute the weight of an arbitrary $m$-bit string is $O(m)$. Since a full adder can be implemented using a constant number of NAND gates, the number of NAND gates required to compute the weight of an $m$-bit string is also $O(m)$.
@@
The inner product requires $O(m)$ OR and AND gates and the weight computation requires $O(m)$ full adders (plus at most $\lceil\log_{2}m\rceil$ AND and OR gates, which we ignore). So the circuit size for computing ${\sf MM}_{n}$ is $O(n)$.
To obtain more concrete estimates, it is convenient to convert the various gate counts into a single unit. Previous works [1, 8] have taken a single NAND gate as the basic unit and translated other gates in terms of this unit. A half-adder can be implemented using 5 NAND gates, while a full adder can be implemented using 9 NAND gates. In [1, 8], a XOR gate was taken to be 2.5 units and an AND gate was taken to be 1.5 units. Between the papers [1] and [8] there is a difference in the number of units required for a flip-flop:  [1] takes a flip-flop to be 8 units, while [8] takes a flip-flop to be 12 units. In Table 2, we provide estimates of the circuit sizes of $\mathcal{S}(L,m)$ for the values of $(L,m)$ in Table 1. These estimates consider a flip-flop to be 8 units.
@@
We note that it is possible to ramp up security at the cost of a reasonable increase in gate count. For example, from Table 1, at the 128-bit security level, our proposal has $L=257$ and $m=59$, resulting in linear bias $2^{-60}$ and the fast algebraic attack requiring more than $2^{130.12}$ operations. If we increase $m$ from 59 to 63, then the linear bias drops to $2^{-64}$, and the fast algebraic attack now requires more than $2^{135.83}$ operations. The gate estimate for $\mathcal{S}(257,63)$, i.e. the stream cipher with $L=257$ and $m=63$, is 2818.5 gates (the LFSR requires 2056 gates, and $f_{127}$ requires 762.5 gates). So the security increases by about 5 bits at the cost of an increase of only 47 gates in the circuit size.

### v2

@@ 4 Concrete Proposals
Given $\kappa$, we obtained representative values of $L$ and $m$. The procedure we followed is to choose the value of $L$ to be the first prime number greater than $2\kappa$, and then for the chosen value of $L$, choose $m$ to be the least integer such that $\beta,\gamma>2^{\kappa}$. This ensures security against algebraic attacks considered above. Next, using $\kappa$ and the corresponding values of $L$ and $m$, we computed the maximum value of $B$ such that $B\leq\kappa$, $\alpha>2^{B}$, and (9), (10) and (11) hold. In each of the cases that we considered, it turns out that this maximum value of $B$ is in fact $\kappa$. So for the chosen values of $L$ and $m$, the corresponding $\mathcal{S}(L,m)$ ensures $\kappa$-bit security against the above analysed correlation attacks even when the adversary has access to $2^{\kappa}$ keystream bits. We note, however, that generating $2^{\kappa}$ keystream bits from a single key and IV pair is meaningless from a practical point of view. Instead, we set $B=64$, i.e. from a single key and IV pair at most $2^{64}$ keystream bits are to be generated. Table 1 shows the values of $\kappa$, $L$, $m$ and other parameters of the filtering function. We note the following points regarding the entries in Table 1.
@@
    The number of variables of the filtering function $f_{2m+1}$ is $12m+$ 1 (and not $m$).
@@
    The feedback polynomial for the LFSR has to be a primitive polynomial of degree $L$. The above analysis of correlation attacks shows that for the obtained values of $L$ and $m$, it does not matter whether the polynomial is sparse, or whether it has a sparse multiple. So from an efficiency point of view, one may choose a low weight primitive polynomial. Examples of primitive polynomials for the values of $L$ in Table 1 are shown in Table 2.
$L$ prim poly
163 $x^{163}\oplus x^{7}\oplus x^{6}\oplus x^{3}\oplus 1$
257 $x^{257}\oplus x^{7}\oplus x^{5}\oplus x^{4}\oplus x^{3}\oplus x^{2}\oplus 1$
331 $x^{331}\oplus x^{7}\oplus x^{6}\oplus x^{5}\oplus x^{4}\oplus x^{2}\oplus 1$
389 $x^{389}\oplus x^{7}\oplus x^{6}\oplus x^{3}\oplus x^{2}\oplus x\oplus 1$
449 $x^{449}\oplus x^{9}\oplus x^{6}\oplus x^{5}\oplus x^{4}\oplus x^{3}\oplus x^{2}\oplus x\oplus 1$
521 $x^{521}\oplus x^{9}\oplus x^{6}\oplus x^{5}\oplus x^{3}\oplus x\oplus 1$
Table 2: Examples of primitive polynomials for the values of $L$ shown in Table 1.
@@ 5 Efficiency of Computing ${\sf MM}_{n}$
In Section 4, we proposed using $f_{2m+1}$ (which is either ${\sf MM}_{2m+1}$ or $1\oplus{\sf MM}_{2m+1}$) as the filtering function in the nonlinear filter model. Further, Table 1 provides specific suggestions of values of $m$ for achieving different security levels. In this section, we consider the complexity of implementing ${\sf MM}_{2m+1}$. Since ${\sf MM}_{2m+1}$ is constructed from ${\sf MM}_{2m}$ using one XOR gate, we need to consider the complexity of implementing ${\sf MM}_{2m}$.
@@
A half-adder takes two input bits and outputs two bits which represent sum of the two input bits. A full adder takes three input bits and outputs two bits which represent the sum of the three input bits. We estimate the numbers of half and full adders that are required for computing the weight of an $m$-bit string $\mathbf{x}$. The algorithm for computing weight that we use is from [6]. If $m=1$, then no adders are required, if $m=2$, a half adder computes the weight, and if $m=3$, a full adder computes the weight. For $m>3$, write $m=m_{1}+m_{2}+1$, where $m_{1}+1$ is the highest power of two that is at most $m$. The algorithm computes the weight of the first $m_{1}$ bits of $\mathbf{x}$, the weight of the next $m_{2}$ bits of $\mathbf{x}$, and then adds these two weights together together with the last bit of $\mathbf{x}$.
Using the above algorithm, it is easy to show that the computation of the weight of an $m$-bit string, when $m=2^{r}-1$ with $r\geq 1$, requires $12^{r}-r-$ 1 full adders. From this it easily follows that the number of full adders required to compute the weight of an $m$-bit string for arbitrary $m$ is $O(m)$. Since a full adder can be implemented using a constant number of NAND gates, the number of NAND gates required to compute the weight of an $m$-bit string is also $O(m)$.
@@
The inner product requires $O(m)$ OR and AND gates and the weight computation requires $O(m)$ full adders. So the circuit size for computing ${\sf MM}_{n}$ is $O(n)$.
Next we consider concrete estimates. For such estimates, we ignore the at most $\lceil\log_{2}m\rceil$ AND and OR gates required for computing the threshold function from the weight, and also the gates required to compute the next bit of the LFSR. Similar assumptions were made in [1] to obtain the gate count estimate for Grain-128a. To obtain concrete estimates, it is convenient to convert the various gate counts into a single unit. Previous works [1, 8] have taken a single NAND gate as the basic unit and translated other gates in terms of this unit. A half-adder can be implemented using 5 NAND gates, while a full adder can be implemented using 9 NAND gates. In [1, 8], a XOR gate was taken to be 2.5 units and an AND gate was taken to be 1.5 units. Between the papers [1] and [8] there is a difference in the number of units required for a flip-flop:  [1] takes a flip-flop to be 8 units, while [8] takes a flip-flop to be 12 units. In Table 3, we provide estimates of the circuit sizes of $\mathcal{S}(L,m)$ for the values of $(L,m)$ in Table 1. These estimates consider a flip-flop to be 8 units.
@@
We note that it is possible to ramp up security at the cost of a reasonable increase in gate count. For example, from Table 1, at the 128-bit security level, our proposal has $L=257$ and $m=59$, resulting in linear bias equal to $2^{-60}$ and the fast algebraic attack requiring more than $2^{130.12}$ operations. If we increase $m$ from 59 to 63, then the linear bias drops to $2^{-64}$, and the fast algebraic attack now requires more than $2^{135.83}$ operations. The gate estimate for $\mathcal{S}(257,63)$, i.e. the stream cipher with $L=257$ and $m=63$, is 2818.5 gates (the LFSR requires 2056 gates, and $f_{127}$ requires 762.5 gates). So the security increases by about 5 bits at the cost of an increase of only 47 gates in the circuit size.